OpenStudy (anonymous):

write each equation in standard form using integers. y=2x+5 y-4=-2(x-3)

OpenStudy (anonymous):

@Ednac

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

So, the standard form of the equation is: \(\large\color{black}{ Ax+yB=C }\) where A, B and C are just constants (numbers, not variables).

OpenStudy (anonymous):

ok i understand so far

OpenStudy (solomonzelman):

I mean, \(\large\color{black}{Ax+By=C }\) sorry.

OpenStudy (solomonzelman):

Okay, so you have to rearrange your equations, so that y and x are on one side and y, and the number is on the other.

OpenStudy (anonymous):

its all good. i just forgot how to change it to standard form

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

I'll do the first one for you, and you do the second, okay?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

My connection snapped, apologize.

OpenStudy (solomonzelman):

\(\large\color{black}{ y=2x+5 }\) subtract 2x from both sides, \(\large\color{black}{ y\color{red}{-2x}=2x+5\color{red}{-2x} }\) 2x on the right cancels. \(\large\color{black}{ y-2x=5 }\) then the Xs have to be before the Ys. \(\large\color{black}{ -2x+y=5 }\) I switched X and Y, I am done.

OpenStudy (anonymous):

ok let me try the second one

OpenStudy (solomonzelman):

sure. go ahead;)

OpenStudy (solomonzelman):

if your stuck, I'll give you a step.

OpenStudy (anonymous):

ok can you help me with the first step

OpenStudy (solomonzelman):

Expand the right side.

OpenStudy (anonymous):

subtract 4 from both sides?

OpenStudy (solomonzelman):

\(\large\color{black}{ y-4=-2(x-3) }\) (it's hard to keep looking back up ) ``` and you can use my thingy with codes: just copy the line below, \(\large\color{black}{ y-4=-2(x-3) }\) ``` And Wait, have you expanded the right side? I mean, you can start from adding 4 from both sides too if you would like to, but you would either way have to expand the \(\large\color{black}{-2(x-3) }\)

OpenStudy (solomonzelman):

having trouble with something, what is it?

OpenStudy (anonymous):

its just confusing to me I'm just going over what you said to do

OpenStudy (solomonzelman):

well, you can ignore the part with codes, if you want.

OpenStudy (solomonzelman):

I re-posted the equation not to look back. Then I gave you the codes to use in a gray box. And after this, I said that you can start from either adding 4 to both sides, or expanding the -2(x-3).

OpenStudy (anonymous):

y=-2(x-3)+4

OpenStudy (anonymous):

is this what it should look like so far

OpenStudy (solomonzelman):

yes, you are going the correct direction, then expand the -2(x-3) (thats on the right hand side )

OpenStudy (anonymous):

i feel stupid asking but how to i expand the right side

OpenStudy (solomonzelman):

Okay, multiply -2 by x, and by -3.

OpenStudy (solomonzelman):

So what will \(\large\color{black}{ -2(x-3) }\) be equal to?